Two times the age of Aaron is 3 more than the age of King. In five years time, the sum of their ages will be 40. Find their present ages.

Answer by ikleyn(52802)

.
Two times the age of Aaron is 3 more than the age of King. In five years time, the sum of their ages will be 40. Find their present ages.
~~~~~~~~~~~~~~~~~~~~~~~~~~~

2A = K + 3            (1)   ("Two times the age of Aaron is 3 more than the age of King")
(A+5) + (K+5) = 40.   (2)   ("In five years time, the sum of their ages will be 40")

Simplify:

2A = K + 3,           (3)
A + K = 30.           (4)

From (3), express  K = 2A - 3  and substitute it into (4). You will get

A + (2A-3) = 30.

At this point I will leave you. Please complete the solution on your own.

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problem for the day of April, 1
    - OVERVIEW of lessons on age problems
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

It contains a lot of other word problems and many other interesting and useful things.